What should be the charge on a sphere of radius 4cm , so that when it is brought in contact with another sphere of radius 2cm carrying charge of 10 micro coulomb , there is no transfer of charge from one sphere to other? 

Grade:12

2 Answers

Manan Kalavadia
2 years ago
Their both potential must be same, so the answer should be, Kq1/r1 = Kq2/r2. on solving you will get 20 micro coloumb.
Manan Kalavadia
2 years ago
Their both potential must be the same, so the answer should be, Kq1/r1 = Kq2/r2. on solving you will get 20 micro coulomb.